Im not sure if this is the correct lesson. \
Directions: Solve the equation. Write solution(s) as intergers, if possible. Otherwise, write as radical expressions
I dont understand what it's asking me.
If you could explain
1. 5 x squared =500


2. x squared + 3 =12



3. x squared + 49 = 0



4. 2x squared - 5 = 27

Answer by stanbon(75887)

1. 5 x squared =500
-----
5x^2 = 500
x^2 = 100
x = +10 or x = -10
===========================
2. x squared + 3 =12
x^2 = 9
x = +3 or x = -3
==========================
3. x squared + 49 = 0
x^2 + 49 = 0
x = 7i or x = -7i
===============
4. 2x squared - 5 = 27
2x^2 = 32
x^2 = 16
x = +4 or x = -4
-------------------------------
Cheers,
Stan H.
==================